Assuming that workers' salaries in your company are uniformly distributed between $20,000 and $50,000 per year, calculate the average salary in your company.
Answer #1

here for uniform distribution: parameter a =20000 and b=50000 ; therefore

average salary in your company =(a+b)/2 =(20000+50000)/2 =35000